# Unravelling the Mechanisms of Manipulating Numbers in Language Models

### v1

@@ 3.3 Different Layers use Interchangeable Numeric Representations
Having found the remarkable degree to which the embeddings for numbers align across different models and the noteworthy universality of sinusoidal representations in varied contexts, we turn to further analyses assessing the extent to which these structures are utilized by the models. First, we address whether the same sinusoidal structure is conserved across a model’s computations.
To that end, we employ the probes we developed in Section 3.2 trained separately for each model layer, $L_{i}$, and measure the accuracy of each probe on representations derived for every other layer $L_{j}\neq L_{i}$. This approach, while informative, comes with the caveat that probes trained for a given layer $L_{i}$ might pick up on idiosyncrasies inherent to a specific layer — i.e., the probes might not disentangle what is specific to numbers as opposed to what is specific to a layer. To address this point, we also fit probes using all but one layer ($L_{1},L_{2},\dots L_{i-1},~L_{i+1},\dots L_{n}$), and evaluate the performance on representations from the held out layer $L_{i}$. As previously, we hold out a subset of 100 numbers for assessing generalization in validation and test conditions.
@@
Figure 5: Probes accuracy on activations from unseen layers.
Results of probes’ cross-layer generalization for Llama 1B are displayed in Figure 5, with largely consistent results for other models in Appendix A.2. We observe that probes fitted on a concrete layer’s representations (Figure 5a) generalize outstandingly well to close-by layers, with the exception of the first and the last layer. This trend reveals that language models operate mostly in a consistent representation of numbers that is universal across their computations, undergoing only minor shifts across layers. This trend is also corroborated by multi-layer probes evaluated in heldout-layer fashion, reaching an accuracy between 95 and 100% across all intermediate layers and three models of diverse sizes and families. This result has practical implications for future work in numeric interpretability — showing that we can train universal, yet highly accurate probes for intermediate layers of a broad set of language models [^4] We will release our training scripts, together with reproducibility guidelines for training universal probes, in a final version of this paper.
This is visualized in Figure 6, showing that intermediate layers in Llama 1B lead to probes with very few weights with values $10^{-5}$.
Nevertheless, we note that this trend is pertinent across different models to a different extent: OLMo 7B and Phi 15B learn more distributed features also across the intermediate layers — which, in turn, also leads to a better generalization of universal probes (Figure 5b), but still causing an outstanding drop in held-out accuracy compared to other layers. Finally, we note that the models’ sparsity profile does not relate to reported performances on arithmetic tasks Kadlčík et al. (2025) and it also does not necessarily determine the accuracy of probes trained specifically for particular layer(s) (Figure 3).
Figure 7: Accuracy of previous tokens’ recovery for different probes trained on embeddings.
Figure 8: Accuracy of extracting input numbers from each component of Transformer block.
Finally, we explore the origin of the high consistency of representations across layers within the internal mechanism that each transformer layer implements. In Figure 8, we visualize the accuracy of probing input representations from each of the components present in transformer layers (Llama 3.2 1B for brevity). We find the accurate numeric representation is scattered across different components in the attention mechanism, with the attention output projection largely violating the sinusoidal representation, which is then, to a large extent, reconstructed in the subsequent fully-connected block. Nevertheless, the consistency of layers’ output representation is primarily maintained by residual streams across layers.
To summarize, we find that the same probe can recover the numeric information for representations pooled from different layers of the models we study. Added to our previous observations in Sections 3.1 and 3.2, we can therefore stress that models learn input embeddings for numbers that are strikingly similar, regardless of which model they come from. These embeddings are processed into sinusoidal representations that are systematic and consistent regardless of context or hidden layer, with the source of this consistency originating primarily in the residual stream across layers.
4 Mechanisms of Number Manipulation
We have established that models converge to sinusoidal number representations universally. This naturally begs the question of how those sinusoidal representations are used.
@@ 4.1 Multi-token Numbers
Our first point of order concerns digits beyond the range of what can be represented in recent LLMs with a single token. To see whether LMs are able to represent even the values of multi-token numbers systematically and accurately, we probe the values of all parts of multi-token numbers from a single representation corresponding to the last numeric token. The accuracy of this experiment uncovers the extent to which the models implement a systematic algorithm of superposing multi-token numbers in a single representation.
@@
For this experiment, we adapt the methodology utilizing natural-language contexts, with a substitution of original numeric values with values spanning two to six numeric tokens, i.e. in a numeric range between $10^{3}-10^{18}$. We use a subset of BBC data that was published after the knowledge cutoff or release date of all the models under consideration (Li et al., 2024). [^5] Retrieved from RealTimeData/bbc_news_alltime
@@
To be able to contextualize the magnitude of superposition and thus, disentangle a general model mechanism from the mechanism specific for numbers, we also probe representations of natural-language tokens using TunedLens (Belrose et al., 2025). TunedLens is an early-exit interpretability method which aims to explain the contents of intermediate representations by learning to translate them into logits. In practice, for each layer $L_{i}$, TunedLens involves distilling the computations done in the subsequent layers $L_{i+1},\dots L_{n}$ into a simple affine transformation $\mathbf{W}_{i}$, such that applying this transformation $\mathbf{W}_{i}$ to a hidden state at $L_{i}$ followed by the unembedding matrix closely matches the logits that the LLM would eventually produce. Here, to maximize comparability and align the number and distribution of target categories, we narrow down the space of probed tokens to 1000 tokens (matching the number probes) ranked as 2000–3000thmost-common tokens in our dataset.
@@
Results for Llama 1B and different probing methods are presented in Figure 8. Results of sin probes show that multi-token numbers are indeed systematically and highly accurately superposed in the latest numeric representation – reaching an accuracy of 99% for the immediately-preceding number piece (offset -1). However, the accuracy and/or systematicity of the superposition mechanism quickly drops for numbers longer than three tokens (i.e. $\geq 10^{9}$), reaching close to zero for the 5th preceding token.
The results of natural-language token probing using TunedLens reveal that superpositioning representations of previous tokens is not a mechanism specific only to numbers – albeit this mechanism seems more prevalent and accurate for numeric tokens. Aiming to maximise comparability, we also report a comparison of the previous-token recovery using TunedLens for numeric tokens — in exactly the same configuration as for the natural-language tokens. This comparison shows that TunedLens is highly inefficient for recovering numeric tokens compared to sinusoidal probes.
4.2 Error Tracking
Establishing that LLMs use universal, systematic and interchangeable representations of numbers, we aim to explore whether we can build upon this knowledge in tracing the origins of models’ outputs in arithmetic reasoning. First, we assess the predictive power of models’ internal representations towards true results in prompts requiring addition, subtraction, multiplication and division.
@@ 4.2 Error Tracking
Figure 9: Probing accuracy in predicting models’ outputs for different prompted operations, for cases where the model predicts a correct (blue) and incorrect response (red).
Results of predicting the Llama 3.2 3B’s expected outputs for different operations are displayed in Figure 9. We can see a large distinction between the cases where the model predicts a correct and incorrect response – suggesting that the extent to which the model maintains the sinusoidal representation may determine the accuracy of the output. We observe that for addition and subtraction, probes can, with close to 100% accuracy, identify the retrieved result already from the model internals. Based on qualitative assessments of output embeddings (subsection B.1), we hypothesise that lower reliability in other operations (multiplication and division) may be caused by the models’ divergence from sinusoidal representation employed by the input/output embeddings, accompanied also by lower accuracy of models overall (Table 2).
@@
((d))
The results in Figure 10 show that across all operations, models tend to incrementally reduce the error towards the true answer value, gradually refining across layers. These results also suggest that there are particular layers responsible for an increase of errors in the model’s internal computation.
((a))
((b))
((c))
((d))
Figure 11: Error aggregation per layer: Relative ratio of cases where each of the layers in Llama 3.2 3B breaks the correct result from the previous layer. We find that removing the three layers with the highest error aggregation in division (layers 5, 9 and 11) leads to 27–64% error reduction.
Figure 11 corroborates this hypothesis — showing that particular layers break the correct result probed from the previous layer in large proportions of all predictions. For instance, in division, the 5th layer is responsible for breaking the correct result recoverable from the previous layer in over 10% of all cases. Our further analyses, overviewed in Table 2, show that according to the probed representations, the models often achieve a correct result internally, even though it does not surface: this is the case of 56.8% of all surfaced (prediction) errors in subtraction, 26.3% in multiplication, and as much as 94.4% of all errors in division.
We further support the claimed responsibility of specific layers on the resulting accuracy by simply removing the suspected layers from the model. Such a naive alternation becomes viable with the uncovered universality of representations across layers (subsection 3.3). Specifically, we try to remove one of the three layers with the highest error aggregation, separately in multiplication and in division (having a potential for improvement of accuracy). We find that this brings performance benefit in four out of six cases – in one case (layer 4) in multiplication, causing a 26% error reduction (from 90.38% to 92.91% and in all cases in division with error reduction between 27–64%.
In summary, factoring in the knowledge of the sinusoidal nature of number representations allowed us to understand and assess the lower bounds of accuracy of multi-token number representations. On a case study of Llama 3.2 3B, we further showed how more accurate probes allow pinpointing – and eliminating – the sources of model errors.
@@ 5 Conclusions
Building upon the recent advancements in probing techniques, this paper builds a fundamental understanding of how LLMs internally represent and manipulate numeric values. We provide evidence that diverse LLMs learn and employ mutually-interchangeable, consistent sinusoidal representations across their internal layers, maintained largely by residual streams, but distinct in the embedding representation and internal activations in terms of sparsity. We show how probing techniques respecting representational properties of numbers open up new possibilities for tracking the causes of errors in models’ internal computation, attributing large portions of errors to numeric interventions of concrete layers. Towards the fast-growing field of interpretability research, our work contributes by evidencing the importance of natural-language probes in training better-generalising probes or the superior quality of specialised probes compared to widely-used methods such as linear or TunedLens probes. We hope that our work clearly outlines the potential of more accurate probes in robustness and faithfulness assessments of existing and future reasoning models.
@@ Acknowledgments
This project has received funding from the European Union’s (EU) Horizon Europe research and innovation programme under Grant agreement No. 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10052546). The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the EU.
Appendix A Experimental details and supplementary results
@@ A.1 Selecting the optimal number of Fourier components for comparison
In Section 3.1, we discuss selecting the optimal number of Fourier components. As can be assessed from LABEL:fig:fft_iou:example, from a certain frequency rank, Fourier decompositions resort to noise. Thus, our objective is to find the “cutoff” rank ($k$) that disentangles oscillating information from the non-oscillating, presumably non-numeric information. This means finding an optimal, yet non-trivial (e.g. $\geq 10$) cutoff for a number of components. In practice, we evaluate for each value of $k$ what the minimum IoU agreement score across all models amounts to for numbers and random overlapping word-pieces. As displayed in Figure 12, we find such an optimal cutoff at $k=63$ as the highest value of $k$ that leads to perfect agreement across models.
A more thorough assessment of Figure 12 suggests a few interesting trends. Random word pieces also tend to favor a handful of Fourier basis components; which we conjecture is due to the sampling mechanism. By selecting overlapping pieces, our pieces must be frequent enough to be present in multiple distinct tokenizers, which in turns shapes the type of linguistic units represented in this random sample. Secondly, we also observe a ‘cross-over’ point around $k\textsection\approx 250$, after which we find greater agreement in random pieces than numbers. Yet, we note that (i) a significant proportion of the mass is concentrated in a few frequency components for numbers, whereas random pieces lead to much more uniform distributions across frequeny components; and (ii) the results in Figure 12 still allow us to establish that the Fourier profile of number pieces is clearly dsitinct from what we observe for any other overlapping set of word pieces.
A.2 Generalization of probes to unseen layers
Figure 13 shows how probes fitted on each layer generalize to all other layers in a model. Strong cross-layer generalization indicates a high consistency of representations throughout the model.

### v2

@@ 3.3 Different Layers use Interchangeable Numeric Representations
Having found the remarkable degree to which the embeddings for numbers align across different models and the noteworthy universality of sinusoidal representations in varied contexts, as a final step, we turn to assess whether the models utilize the identical representation throughout its computation. To investigate that, we employ the probes from Section˜3.2 trained separately for each separate model layer ($L_{i}$) and measure the accuracy of each probe on representations derived for every other layer $L_{j}\neq L_{i}$.
While informative, this approach still comes with the caveat that probes trained for a given layer $L_{i}$ may pick up on idiosyncrasies inherent to a specific layer — i.e., the probes might not disentangle what is specific to numbers as opposed to what is specific to $L_{i}$. To address this point, we also fit probes using all but one layer ($L_{1},L_{2},\dots L_{i-1},~L_{i+1},\dots L_{n}$), and evaluate performance on representations from the held out layer $L_{i}$. As before, we hold out 100 numbers for assessing generalization in validation and test conditions.
@@
Figure 5: Transferability of representations: Accuracy of param-sin probe evaluated on unseen layers
Results of probes’ cross-layer generalization for Llama 1B are displayed in Figure˜5. We observe that probes fitted on a concrete layer’s representations (LABEL:fig:one-v-one) generalize outstandingly well to both close-by and more distant layers, with the exception of the first and the last layer. This trend reveals that language models operate in a consistent representation of numbers that is universal throughout their computation, undergoing only minor shifts across layers. This trend is also corroborated by multi-layer probes evaluated in heldout-layer fashion (LABEL:fig:all-v-one), reaching a close-to 100% accuracy across all intermediate layers and six models of diverse sizes and families. This result has practical implications for future work in numeric interpretability — showing that we can train universal, yet highly accurate probes for intermediate layers of a broad set of language models.
To summarize, we find that the same key, i.e. probe can be used to interpret numeric information at different layers of the same language model. Added to our previous observations in Sections˜3.1 and 3.2, we can therefore stress that models learn input embeddings for numbers that are strikingly similar, regardless of which model they come from. These embeddings are processed into sinusoidal representations that are systematic and consistent regardless of context or hidden layer, with the source of this consistency maintained primarily by the residual stream across layers (cf. Section˜A.7).
4 How to Study Numbers in LLMs
We have established that models universally converge to sinusoidal number representations. This characteristic is not shared with other word pieces. This questions the appropriateness of existing interpretability tools when it comes to numbers and other ordinal data types.
Figure 6: Accuracy of extracting pieces of multi-token numbers from the last token representations.
@@ 4.1 Multi-token Numbers
As a case study for observing how the choice of interpretability toolset affects our findings, we analyse whether our conclusions about the remarkable accuracy of numeric representations also hold for large, multi-token numbers.
@@
First, we use our sinusoidal probes to retrieve the values of all parts of multi-token numbers from a single, i.e. last token’s representation. For this experiment, we adapt the methodology utilizing natural-language contexts (see Figure˜4), with a substitution of original numeric values with values spanning two to six numeric tokens, i.e. in a numeric range between $10^{3}-10^{18}$. We use a subset of BBC data that was published after the cutoff or release date of our LLMs (Li et al., 2024).
@@
As a point of comparison, we consider the general-purpose TunedLens tool (Belrose et al., 2025), an early-exit interpretability method which aims to explain the contents of intermediate representations by learning to translate them into logits. For each layer $L_{i}$, TunedLens distills the computations done in the subsequent layers $L_{i+1},\dots L_{n}$ into a simple affine transformation $\mathbf{W}_{i}$, such that applying this transformation $\mathbf{W}_{i}$ to a hidden state at $L_{i}$ followed by the unembedding matrix closely matches the logits that the LLM would eventually produce. TunedLens was applied to number representations in previous work (Zhou et al., 2024) and also represents the line of work arguing for a linear representations of numeric values in LLMs (Zhu et al., 2025). To maximize comparability and align the number and distribution of target categories, we narrow down the space of probed tokens to 1,000 tokens ranked as 2,000–3,000th most-common tokens in our dataset (matching the number of classes and their distribution for our number probes). Crucially, we also report previous-token recovery rates using TunedLens for number tokens, using the same configuration as for natural-language tokens.
@@
Results for Llama 1B and different probing methods are presented in Figure˜6. Results of sin probes show that multi-token numbers are indeed systematically and highly accurately superposed in the latest numeric representation – reaching an accuracy of 99% for the immediately-preceding number piece (offset -1). However, the accuracy and/or systematicity of the superposition mechanism quickly drops for numbers longer than three tokens (i.e. $\geq 10^{9}$), reaching close to zero for the 5th preceding token.
We notice that TunedLens gives a conflicting account. While it shows that previous non-number tokens are also superposed, i.e. can be decoded as far as five time-steps later with non-zero probability, it suggests that this is not the case for number tokens – these cannot be retrieved past two time-steps, and even then, only with 50% accuracy. However, a comparison with the sin probe shows that this trend is merely a result of TunedLens’s inefficiency for numeric tokens; estimates derived from it provide an inaccurate representation of LLMs’ accuracy, underestimating the amount of contained information and overlooking the superposition mechanism uncovered by a more suitable probe.
Llama 3.2 3B   Refer to caption Refer to caption Refer to caption Refer to caption
OLMO 2 7B   Refer to caption Refer to caption Refer to caption Refer to caption
Phi 4 15B   Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Probing accuracy in extracting correct output from internal activations across layers; different probes.
4.2 Output Tracking
A reliable mechanism of numeric manipulations necessitates an accurate representation of inputs (evidenced in §3), but also the ability to construct systematic and accurate representations of outputs. Without the specialized tools to analyze numeric representations, prior work describes LLMs’ arithmetical mechanisms through qualitative assessments limited to the outputs of addition Kantamneni and Tegmark (2025); Zhou et al. (2024); Levy and Geva (2025). Therefore, having found the sinusoidal probes to be highly suitable for describing the representations of input numbers, we set off to investigate whether such probes can also better capture the mechanics of LLMs’ output representations, compared to other probe types.
In this set of experiments, we prompt the language models with prompts of the form “The result of $x_{1}\otimes x_{2}$ is  [^4] Exact prompt formulations for each operation can be found in the supplementary material under /notebooks/output_probing.  and train probes to predict the correct result of $x_{1}\otimes x_{2}$, where ${+,,×,/}\otimes\in\{$$\times$,/\}.
Figure 7 shows the results for probes from prior work, including linear probes from Zhu et al. (2025); Zhou et al. (2024) and a sin probe proposed and applied by kadlčík2025pretrainedlanguagemodelslearn. We can see that the param-sin probe performs comparable or better than probes applied for numbers in previous work and newly allows the tracking of outputs with over 80% of accuracy in multiplication. The results for addition also push the known lower bounds of precision of the result emerging early in the model, conflicting with a hypothesis of “gradually refining predictions” from Zhou et al. (2024) based on LogitLens, but supporting the interpretation of arithmetics implementation in LLMs by Stolfo et al. (2023) through a combination in late layer(s).
@@ 4.2 Output Tracking
Figure 8: Probing accuracy in extracting correct output from Llama 3.2 3B’s internal activations separates model’s correct (blue) and incorrect (red) predictions.
Figure˜8 further details on a distinction between the cases where the model predicts a correct and incorrect response – suggesting that the extent to which the model maintains the sinusoidal representation may determine its accuracy.
We evidence this by following qualitative and mechanistic evidence. In Section˜A.5, we observe the output embeddings of addition and multiplication – latter coupled with lower model’s accuracy – differ in the precision of the the observed sinusoidality; with the pattern being more prominent in the better-performing, addition embeddings. In Section˜A.8, we further analyze how, according to our probes, is the result refined via gradually minimizing the response error across layers, with particular layers deteriorating this refinement, suggesting a worse fit of the sinusoidal probe and thus, the sinusoidal pattern. We find that simply dropping the layers causing a divergence from the probe’s fitted pattern lead to an accuracy improvement in multiplication and division in 4 out of 6 cases.
Finally, we evidence this point more directly by a mechanistic intervention – steering activations in prediction cases leading to incorrect prediction towards the expected sinusoidal pattern. Concretely, we (i) fit sinusoidal probes at each layer to predict the output of a given arithmetic operation; (ii) optimize a set of 1,000 randomly-initialized ‘patch’ embeddings $\mathbf{e}_{y}$ to maximize the probability that the probes will map them onto the integer value $y$; and (iii) steer activations towards the corresponding patch whenever the model is not producing the expected output. Steering is achieved by interpolating between the activation $\mathbf{h}_{i}$ at layer $L_{i}$ and the patch $\mathbf{e}_{y}$ for the target scalar $y$, $\alpha\mathbf{h}_{i}+(1-\alpha)\mathbf{e}_{y}$. We report results with respect to the optimal value of $\alpha$ per layer, operation and probe.
@@
Figure 9: Error reduction after steering activations on incorrect predictions of multiplication and division. Optimal interpolation factor $\alpha$ per setup.
Figure˜9 displays the absolute difference in performance caused by the steering on previously erroneous cases in multiplication and division for three models. We target probes on upper layers which reach accuracy scores $\geq 10\%$. In short, depending on the layer, we can correct up to $30\%$ of errors for multiplication, and up to $42\%$ for division. Crucially, our probe almost always outperform the probe proposed by kadlčík2025pretrainedlanguagemodelslearn: Across all setups, we only observe four setups where the sin probe outperforms our sin-param probe. [^5] Viz., division with OLMo at L22, L23, L24), multiplication with Phi at L37. The optimal setup per model and operation is always obtained with our probe. This underscores once again the importance of selecting the right explanatory tool, this time in the context of activation steering.
In sum, factoring in the knowledge of the sinusoidal nature of number representations allowed us to understand and assess the lower bounds of accuracy of multi-token number representations. On a case study of Llama 3.2 3B, we further showed how more accurate probes allow pinpointing and eliminating sources of errors. Tools that are best suited to number representations allow us to enhance accuracy and perform more effective interventions.
4.3 Ordinal Data Types
Having established that sinusoidal mechanisms robustly encode numeric values and our specialized probe facilitates targeted interventions, we investigate how this tooling generalizes to ordinal data types – potentially represented in the same systematic fashion as numbers. A particular challenge here is a data scarcity – for instance, as there are only seven days in a week, to warrant held-out evaluation, we can only fit the probes on six examples.
Prior work by Engels et al. (2025) demonstrates that LLMs internally represent temporal data – such as days of the week and months of the year in circular structures. Analytically, a centered circle is parameterized by the equations $x(\theta)=r\sin(\theta+\pi/2)$ and $y(\theta)=r\sin(\theta)$, indicating that these temporal concepts are encoded using sinusoidal features. Assuming this structure, we fit our probe on the embedding representations of these concepts and compare it to a prior probe architecture of kadlčík2025pretrainedlanguagemodelslearn in leave-one-out cross-validation setting. The best held-out validation accuracies of both architectures are reported in Table˜1.
Model sin param-sin
Days of week (n=7)
Olmo 2 1B 4 6
Olmo 2 7B 6 6
Olmo 2 13B 5 6
Llama 3 1B 5 6
Llama 3 3B 7 6
Llama 3 8B 3 6
Months (n=12)
Olmo 2 1B 6 11
Olmo 2 7B 6 11
Olmo 2 13B 5 11
Llama 3 1B 9 11
Llama 3 3B 10 11
Llama 3 8B 6 11
Numerals: ‘one’ to ‘ten’ (n=10)
Olmo 2 1B 3 6
Olmo 2 7B 4 8
Olmo 2 13B 4 8
Llama 3 1B 3 5
Llama 3 3B 4 6
Llama 3 8B 3 9
Table 1: Number of matches achieved by training param-sin and best-performing prior probe (sin) in decoding ordinal data with extremely small training sets from models’ internals. Held-one-out evaluations.
Our results demonstrate high decoding accuracy for ordinal data across all evaluated models. This validates the probe’s utility beyond pure numbers, even under extreme data-scarcity. With our better generalizing architecture, we corroborate and extend the discovery by Engels et al. (2025) of circularity in LLMs’ representations to other ordinal data types.
@@ 5 Conclusions
In this paper, we have shown the number representations in LLMs are exceptional in two distinct but related ways. First, the regularity and systematicity of the sinusoidal structure that all LLMs converge is striking, to the point of being almost universal across all experimental setups we consider. We can view the same frequencies in Fourier being used by all models we survey (§3.1), we can train a probe on one layer and apply to another (§3.3). The one remarkable exception is the discrepancy between natural language contexts, which yield remarkably robust probes that generalize broadly, and synthetic probes derived from templated texts, which do not.
Second, this exceptional sinusoidal regularity requires adapting the way we study number processing in LLMs. Using a general-purpose TunedLens method, applied for numbers in prior work, grossly underestimates how accurately previous timestep information can be retrieved, whereas sinusoidal probes such as the one we propose in eq.˜2 may lead to substantially higher accuracies and contradictory conclusions (§4.1). Finally, we demonstrate that the sinusoidal probes are also well applicable for analyzing the emergence and manipulation of model’s own predictions and that the accuracy of representation in terms of sinusoidality is related to the prediction accuracy (§4.2).
While we focus on numbers, our experiments showcase how important the appropriate interpretability toolset is for accurate interpretation of specialized data types. This naturally questions the common practice of designing general-purpose interpretability and explainability tools, motivating future work towards adapting our practices to the specifics of the type of data we want to explain.
Limitations
Several practical factors limit the scope of our work. We focus our survey on three particular model families; moreover, we have excluded very large models (>30B parameters) due to practical hardware limitations. The rapid pace of modern NLP may entail that the set of models we have picked for this study will no longer be representative of contemporary frontier models in the future.
Our work is also based on English-centric LLMs, but the number systems vary greatly across languages — e.g., Chinese and Japanese often employ myriad-based number systems and frequently use both Chinese and Arabic numerals. The extent to which our conclusions will hold for a Chinese-centric or Japanese-centric LLM remains a topic for future research.
Lastly, we find that all errors in probing temporal concepts (days, months) are exactly in the first position (Monday, January), persisting even when the probed token list is reversed (Sunday, December). We suspect that zero value is a blind spot of the probe due to its parametrization, which needs to be addressed to achieve a perfect decoding rate.
@@ Acknowledgments
This project has received funding from the European Union’s (EU) Horizon Europe research and innovation programme under Grant agreement No. 101070350. The work has also received funding from the Digital Europe Programme under grant agreement No 101195233 (OpenEuroLLM). The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the EU.
Appendix A Supplementary results
@@ A.1 Selecting the optimal number of Fourier components for comparison
In Section˜3.1, we discuss selecting the optimal number of Fourier components. As can be assessed from LABEL:fig:fft_iou:example, from a certain frequency rank, Fourier decompositions resort to noise. Thus, our objective is to find the “cutoff” rank ($k$) that disentangles oscillating information from the non-oscillating, presumably non-numeric information. This means finding an optimal, yet non-trivial (e.g. $\geq 10$) cutoff for a number of components. In practice, we evaluate for each value of $k$ what the minimum IoU agreement score across all models amounts to for numbers and random overlapping word-pieces. As displayed in Figure˜10, we find such an optimal cutoff at $k=63$ as the highest value of $k$ that leads to perfect agreement across models.
A more thorough assessment of Figure˜10 suggests a few interesting trends. Random word pieces also tend to favor a handful of Fourier basis components; which we conjecture is due to the sampling mechanism. By selecting overlapping pieces, our pieces must be frequent enough to be present in multiple distinct tokenizers, which in turns shapes the type of linguistic units represented in this random sample. Secondly, we also observe a ‘cross-over’ point around $k\approx 250$, after which we find greater agreement in random pieces than numbers. Yet, we note that (i) a significant proportion of the mass is concentrated in a few frequency components for numbers, whereas random pieces lead to much more uniform distributions across frequency components (see Figure˜11); and (ii) the results in Figure˜10 still allow us to establish that the Fourier profile of number pieces is clearly distinct from what we observe for any other overlapping set of word pieces.
Figure 11: Contributions of Fourier base frequencies for non-number pieces (Llama-3 1B, provided as an example)
A.2 Natural-language contexts for representations extraction
Contrary to prior work, we use the contexts of natural language to extract the representations of numbers. We argue that such setting better corresponds to the practical applicability of probes – in monitoring models’ behavior and errors, or in tracking the faithfulness of reasoning chains.
We choose the contexts with medium to high co-dependency on numerical contents covering four distinct domains and eight different datasets. The datasets categorized into these domains are presented in Table 3 and total in around 700 thousands natural-language contexts.
A.3 Generalization of probes to unseen layers
Figure˜12 shows how probes fitted on each layer generalize to all other layers in a model, using synthetic mathematical contexts. Strong cross-layer generalization indicates a high consistency of representations throughout the model.